OpenStudy (anonymous):

A spacecraft's radio signals traveled for four hours at the speed of light to reach Earth. How far away was the spacecraft? Give your answer in kilometers, using powers of ten notation.

OpenStudy (anonymous):

Lets begin with the speed of light: c=3x10^8 m/s and t=4 hrs=4*3600 s=14.400 s=1.44x10^4 s d=c*t d=............... find it yourself :D

OpenStudy (anonymous):

I have no idea what any of that stands for.

OpenStudy (anonymous):

Lol c=speed of light d=distance t=time

OpenStudy (anonymous):

I need it dumbed down. Please help me, I have no idea what I'm doing. Sorry :/

OpenStudy (anonymous):

Would the distance be in meters?

OpenStudy (anonymous):

yes d in meters.

OpenStudy (anonymous):

okay, thank you for that. it explains it much better but i need the final answer in km. how do i do that?

OpenStudy (shane_b):

First convert hours to seconds: \[4\cancel{hrs}*\frac{60\cancel{min}}{1\cancel{hrs}}*\frac{60secs}{1\cancel{min}}=14,400secs\]The speed of light is 300,000 kilometers/sec so now you just multiply that to get the total distance traveled: \[\frac{300,000km}{\cancel{secs}}*14,400\cancel{secs}=4.32 x 10^9 km\]

OpenStudy (anonymous):

Oh my god, thank you. I'll keep that for future reference!